a small object is released from rest and falls 1.00 * 10^2 feet near the surface of the earth. neglect air resistance. how long will it take to fall through the 1.00 * 10^2 feet mentioned?

a) 4.5 s

b) 10 s

c) 2.49 s

d) 3.12 s

e) 6.25 s
Answer #1

ANS c) 2.49 s

Since 100 feet = 30.48 m

Applying equation s = ut + 1/2 at2

30.48 = 0 + 1/2 *(9.8 * t^2)

t^2 = 30.48 / 4.9 = 6.22

t = 2.49 s